OpenStudy (anonymous):

Part 1: Using complete sentences, explain which method you would use to solve the following system of equations and why you chose that method. 4x + 3y = 2 3x + 2y = 1 Part 2: Show your work and provide the solution to the system.

OpenStudy (anonymous):

t's a tossup 321. You can either solve on of the equations for x and use substitution on the other equation. Or you can do whatever multiplication is needed to make the coefficients on both equations the same and treat it as parallel equations. one and one likes the parallel route. I like it too. But he biffed on subtracting the equations. it should be -x = 1 then y = 2 But i'm quicker with the first option. 3x + 2y = 1 2y = 1 - 3x y = (1-3x) / 2 Now sub this new definition of y into the first equation 4x + 3 ((1-3x) / 2) = 2 4x + 3/2 - 9x/2 = 2 4x - 9x/2 = 2 - 3/2 -1/2 (x) = 1/2 x= -1 since we've already shone y = (1-3x) / 2 y = 2 Run the numbers thru the 2 equations to check.